A horizontal spring-mass system oscillates on a frictionless table. If the ratio of the mass to the spring constant is 0.037 kg·m/N, and the maximum speed of the mass was measured to be 4.5 m/s, the maximum extension of the spring will be  cm.


1
Expert's answer
2021-01-14T10:39:39-0500

Apply energy conservation to solve the problem:


"\frac12kx^2=\frac12 mv^2,\\\space\\
x=v\sqrt{\frac mk}=4.5\sqrt{0.037}=0.866\text{ m},"

or 86.6 cm.